# ul. Pasteura 5, 02-093 Warsaw, Poland

### v1

@@ 3.3 Schematic timeline
We have seen how a typical stringy, post-inflationary cosmology can be approximated by a sequence of distinct epochs, each dominated by different components. Each epoch is characterized by a constant equation of state, with relatively short transients (lasting around one e-fold) in between. A rather general case, studied in this paper, is when the evolution is driven by the rolling of the universal volume modulus. Let us now summarize for later convenience the phases arising in this case, in the context of the Large Volume Scenario or any other decaying exponential potential for the volume modulus, sketched in Fig. 1 (see also Table 1 of Apers:2024ffe ).
Figure 1: Schematic timeline of the scenario. Adapted from Cicoli:2023opf .
After inflation, there will be a kination phase, followed by a brief epoch of radiation/matter domination and then a radiation/matter tracker. Once the volume reaches the vacuum, it will give rise to early matter domination by oscillating around the minimum of the potential, and finally reheat the Universe by decaying to radiation. While the different transients between these epochs can partly be described analytically, we will not give all of the details here, and the interested reader is referred to Section 2 of Apers:2024ffe . In any case, we will also make us of numerical solutions to (33) to calculate GW spectra. With these assumptions in mind, a schematic timeline proceeds as follows. Kination begins at a time
$t_{0}=\frac{M_{P}}{\sqrt{3}\Lambda^{2}_{{\rm inf}}},$ (53)
where $\Lambda_{{\rm inf}}$ denotes the scale of inflation, i.e. $\Lambda_{{\rm inf}}=V_{0}^{1/4}e^{-\lambda\frac{\Phi_{0}}{4M_{P}}}$. During this phase, as discussed around Eq. (50), the string tension will evolve as $G\mu(t)\sim 1/t$. Let us now assume the existence of another fluid, with equation of state $P=(\gamma-1)\rho$ and abundance $\Omega_{\gamma}\equiv\varepsilon_{0}\ll 1$. Neglecting the potential, one finds Apers:2024ffe
$x(a)=\frac{1}{\sqrt{1+\varepsilon_{0}\left(\frac{a}{a}\right)^{3(2-\gamma)}}},$ (54)
@@
$t_{\gamma}\simeq t_{0}\varepsilon_{0}^{\frac{1}{\gamma-2}}.$ (55)
@@
$t\simeq t_{0}\varepsilon_{0}^{\frac{4+\gamma\left(\lambda\sqrt{\frac{2}{3}}-2\right)}{4(\gamma-2)}},$ (56)
@@
$t_{t}\simeq t_{0}\varepsilon_{0}^{\frac{1+\frac{\lambda}{2}+\sqrt{\frac{2}{3}}-\frac{2}{3}}{\gamma-2}}.$ (57)
@@
$t_{{\rm reh}}\simeq\frac{1}{\Gamma_{\Phi}},\quad\quad{\rm{with}}\quad\quad\frac{1}{48\pi}\frac{m_{\Phi}^{3}}{M_{P}^{2}}\lesssim\Gamma_{\Phi}\lesssim\frac{(\alpha_{{\rm{loop}}}\mathcal{V})^{2}}{16\pi}\frac{m_{\Phi}^{3}}{M_{P}^{2}}.$ (58)
@@
To close off, let us remark that the above picture is indeed schematic, with the equations being correct up to $\mathcal{O}(1)$ factor and intended to provide simple, analytical approximations to the evolution of the background useful for an intuitive understanding of the physics at play. For the actual computations of GW spectra presented in the next sections, the values of the various background cosmological parameters (scale factor, Hubble rate, string tension…) are computed numerically. In particular, we numerically solve the system (33) from the end of inflation to the beginning of the matter-dominated epoch, when the field starts oscillating around the minimum. From there onwards, it is simpler to assume a standard radiation-matter universe, with a matter component that decays to radiation as described in 3.1.2 and with initial conditions determined by the previous evolution. Finally, this can be patched to the standard cosmological history before the onset of BBN.
@@ 4 The GW background
In this Section we follow Cui:2018rwi and provide analytic estimates for the amplitude and spectral index of the GW spectrum sourced by a network of cosmic superstrings, as a function of the equation of state of the background when the loops are created and when they source the dominant contribution to the spectrum. In Section 4.1 we reproduce the results of Cui:2018rwi , generalizing them to include expressions for the amplitude as a function of $G\mu$, $\alpha$, $P$ and $\chi$ (the latter is defined in Eq. (14)). In Section 4.2, we generalize the computation to the case where the tension varies as a power law in time, as motivated by moduli dynamics in string theory constructions. We observe that this variation can render qualitatively different changes in the GW spectrum. In particular, it induces spectral indices that cannot be reproduced in scenarios with constant tension, and the maximum contribution to the amplitude at a given frequency is sourced at the time of formation of the loop (while with constant tension this occurs at half of its lifetime).
Let us thus compute the spectral shape of the GW background released by the loops (see Sousa:2024ytl for a recent review). First, we assume that the loops with length $\ell$ radiate in discrete harmonics given by
$\tilde{f}_{k}=\frac{2k}{\ell}\,,\qquad k\in\mathbb{N}^{+}$ (59)
@@
[…]
This equation sums over all times $\tilde{t}$ the (redshifted) energy in GWs sourced from loops with length $\ell(\tilde{t})=\frac{2a(\tilde{t})}{fa_{0}}$. The density of the number of loops per unit length is written in terms of the (redshifted) initial loop distribution, taking special care of the variation of the size of the interval from $(t_{i},t_{i}+dt_{i})$ to $(\ell(\tilde{t}),\ell(\tilde{t})+d\ell(\tilde{t}))$. In addition, $E^{(1)}_{\rm{GW}}$ denotes the fraction of energy that is released into the first mode. In general, one writes
$\frac{dE_{\rm{GW}}}{d\tilde{t}}=\sum_{k=1}^{\infty}\frac{dE^{(k)}_{\rm{GW}}}{d\tilde{t}}\,,$ (61)
which in the language of Eq. (18) corresponds to
$\Gamma=\sum_{k=1}^{\infty}\Gamma^{(k)}\equiv\sum_{k=1}^{\infty}\frac{\Gamma}{k^{4/3}\zeta(4/3)}\,,$ (62)
@@
Throughout the rest of the paper, we use $\rho_{c}^{-1}=\frac{8\pi}{3}\frac{G}{H_{0}^{2}}$ and Eqs. (16) and (62) to write the contribution from the mode $k$ to the gravitational wave spectrum as:
[…]
The Heaviside theta functions ensure that the loop formation time cannot be extrapolated before the network is formed, and impose a high frequency cutoff in terms of a minimum length $\ell_{*}$ that loops can reach before other channels become more efficient (see Ghoshal:2023sfa and references therein). Of course, the total spectral shape is given by the sum of these contributions, namely
$\Omega_{\rm{GW}}(f)=\sum_{k=1}^{\infty}\Omega_{\rm{GW}}^{(k)}\,,$ (64)
@@ 4.1 Frequency dependence: constant tension
As such, one of the most interesting features of cosmic strings is that, if they are present, because of the properties of the scaling regime, their associated GW background carries information about the cosmology following their formation. In particular, the frequency dependence of the spectrum will vary depending on the equation of state of the background. In addition, a time-dependent string tension - as would be expected if the volume modulus was rolling down its potential - leads to different spectral dependence and thus provides a smoking-gun signature for scenarios as in Apers:2022cyl ; Conlon:2022pnx ; Apers:2024ffe .
The background behavior of the spectral index is well understood Cui:2018rwi ; Gouttenoire:2019kij , but a similar analysis with a time-dependent tension has only been carried out in a special case Emond:2021vts . Let us first review the case where the string tension is constant and solely focus on the behavior of the first mode; higher modes yield a qualitatively different behavior only in the case of matter domination (the effect of higher modes has been studied analytically in Gouttenoire:2019kij ). To do so, we follow the logic in Cui:2018rwi and generalize their approach to include the dependence on the intercommutation probability and the parameter $\chi$. Let us consider two cosmological phases with
$\rho(t)=\begin{cases}\rho_{\Delta}\left(\frac{a_{\Delta}}{a(t)}\right)^{m}\,,\qquad&t<t_{\Delta}\\ \rho_{\Delta}\left(\frac{a_{\Delta}}{a(t)}\right)^{n}\,,\qquad&t\geq t_{\Delta}\end{cases}\,.$ (65)
In what follows we assume that the scaling regime is maintained throughout an immediate transition between the two phases in favor of an analytic understanding of the situation, but note that these issues have been studied in the past Sousa:2013aaa ; Sousa:2016ggw ; Avgoustidis:2025svu .
@@
$\left(\frac{a(t_{i})}{a(\tilde{t})}\right)^{3}=\left(\frac{a(t_{i})}{a_{\Delta}}\right)^{3}\left(\frac{a_{\Delta}}{a(\tilde{t})}\right)^{3}=\left(\frac{t_{i}}{t_{\Delta}}\right)^{6/m}\left(\frac{t_{\Delta}}{\tilde{t}}\right)^{6/n}\,.$ (66)
The behaviour of the spectrum with respect to $f$ is understood by studying, at a fixed $f$, which is the time $\tilde{t}$ where the loop population sources the largest contribution to the GW amplitude. Notice that the formation time carries information about the length of the loop at time $\tilde{t}$, from (19) we have:
$t_{i}=\frac{1}{\alpha P^{\chi}+\Gamma G\mu}\left(\frac{2k}{f}\frac{a(\tilde{t})}{a_{0}}+\Gamma G\mu\tilde{t}\right)\,,\qquad\frac{dt_{i}}{d\ell(\tilde{t})}=\frac{1}{\alpha P^{\chi}+\Gamma G\mu}.$ (67)
@@
$f_{\Delta}\equiv\frac{2}{\alpha P^{\chi}}\frac{1}{t_{\Delta}z_{\Delta}}\,,\qquad\frac{t_{i}}{t_{\Delta}}\simeq\frac{f_{\Delta}}{f}\left(\frac{\tilde{t}}{t_{\Delta}}\right)^{2/n}+\frac{\Gamma G\mu}{\alpha P^{\chi}}\frac{\tilde{t}}{t_{\Delta}}\,,$ (68)
where here and in what follows we neglect $\Gamma G\mu$ over $\alpha P^{\chi}$, and $z_{\Delta}$ is the redshift at the time $t_{\Delta}$. Notice that keeping the second term in Eq. (68) does not contradict this statement since the relative terms could be of the same order for a choice of times and frequencies. Indeed, it is precisely this regime that turns out to be interesting for loops in cosmologies with constant string tension (and we will see how this need not be the case otherwise).
@@
$h^{2}\Omega_{\rm{GW},1}^{(1)}=A_{\Delta}\left(\frac{f}{f_{\Delta}}\right)^{3-6/m}\int_{t_{F}}^{t_{M}(f)}\frac{d\tilde{t}}{t_{\Delta}}\,\left(\frac{\tilde{t}}{t_{\Delta}}\right)^{2(6/m-4)/n+4/n}\,.$ (69)
$h^{2}\Omega_{\rm{GW},2}^{(1)}=A_{\Delta}\left(\frac{f}{f_{\Delta}}\right)^{-1}\left(\frac{\Gamma G\mu}{\alpha P^{\chi}}\right)^{6/m-4}\int_{t_{M}(f)}^{t_{0}}\frac{d\tilde{t}}{t_{\Delta}}\,\left(\frac{\tilde{t}}{t_{\Delta}}\right)^{6/m-4+4/n}\,.$ (70)
@@
$A_{\Delta}=\frac{8\pi}{3}\frac{\Gamma^{(1)}}{(H_{0}t_{0})^{2}}\frac{\mathcal{F}_{\alpha}}{\alpha P^{2(1-\beta)+\chi}}\left(G\mu_{\Delta}\right)^{2}\,C_{\rm{eff}}\left(\frac{t_{0}}{t_{\Delta}z_{\Delta}^{2}}\right)^{2}\equiv A\left(G\mu_{\Delta}\right)^{2}\,C_{\rm{eff}}\left(\frac{t_{0}}{t_{\Delta}z_{\Delta}^{2}}\right)^{2}\,,$ (71)
where $A$ is universal and we define $G\mu_{\Delta}\equiv G\mu(t_{\Delta})$ anticipating the case where the tension varies. We note here that the effect of varying the intercommutation probability is a uniform boost in the amplitude for $\chi=0$ (as in Ellis:2023tsl ), and an additional contribution to the amplitude and a shift in the characteristic frequency $f_{\Delta}$ for $\chi\neq 0$.
In addition, we observe that the frequency dependence of the spectrum depends on the exponents in Eqs. (69) and (70), and therefore in the equation of state of the fluid governing the expansion of the Universe. For values of interest $(m,n)\in\{3,4,6\}$ the first exponent of $\tilde{t}/t_{\Delta}$ ($p_{1}$) is always larger than $-1$ while the second ($p_{2}$) is larger than $-1$ except in a matter-radiation transition (i.e. at the end of a period of early matter domination) and during matter domination. We will have more to say about such periods in Sec. 5, since the spectrum is sensitive to higher modes, but let us for the moment assume that $p_{1}>-1$ and $p_{2}<-1$. Then, the integral is dominated at a time $t_{M}$ defined as the time at which the two contributions to $t_{i}$ are equal, namely
$\frac{t_{M}}{t_{\Delta}}\equiv\left(\frac{f_{\Delta}}{f}\frac{\alpha P^{\chi}}{\Gamma G\mu}\right)^{n/(n-2)}\,.$ (72)
@@
$h^{2}\Omega_{\rm{GW}}^{(1)}\simeq A_{\Delta}\left(\frac{f}{f_{\Delta}}\right)^{B}\,\left(\frac{\alpha P^{\chi}}{\Gamma G\mu}\right)^{C}\,.$ (73)
Where $B=2(mn-m-3n)/(m(n-2))$ and $C=(12+mn-4m)/(m(n-2))$. There are subtleties to the analysis above that will be important later. Firstly, we must note that we have assumed that $t_{M}(f)<t_{0}$, where $t_{0}$ indicates the onset of a different phase of the Universe (eg: radiation-matter transition), or the age of the Universe today. Otherwise the dominant contribution to the spectrum at this frequency must be computed in the later phase with times $t>t_{0}$ or, in the case of the latest epoch of matter domination, it is just evaluated today. In the latter case:
$h^{2}\Omega_{\rm{GW}}^{(1)}\simeq A_{0}\left(\frac{f}{f_{0}}\right)^{3-6/m}\,.$ (74)
@@
$f<f_{F}\left(\frac{\alpha P^{\chi}}{\Gamma G\mu}\right)^{2/n}2^{1-2/n}\,,$ (75)
@@
At larger frequencies one simply evaluates (70) at $t_{F}$, finding
$h^{2}\Omega_{\rm{GW}}^{(1)}\simeq A_{F}\left(\frac{f}{f_{K}}\right)^{-1}\left(\frac{\alpha P^{\chi}}{\Gamma G\mu}\right)^{6/m-4}\,,$ (76)
@@
$n=3$ $-1$ $-1$ $1$
$n=4$ $-1$ 00 $1$
$n=6$ $-1/2$ $1/4$ $1$
Table 1: Expected spectral index of the GW spectrum sourced by the first mode of loops that are created when the background energy density redshifts like $\rho\sim a^{-m}$ and radiate GWs with $\rho\sim a^{-n}$, according to Eq. (73). Negative spectral indices are altered by the effect of higher modes to $-1/3$.

### v2

@@ 3.3 Schematic timeline
We have seen how a typical stringy, post-inflationary cosmology can be approximated by a sequence of distinct epochs, each dominated by different components. Each epoch is characterized by a constant equation of state, with relatively short transients (lasting around one e-fold) in between. A rather general case, studied in this paper, is when the evolution is driven by the rolling of the universal volume modulus. Let us now summarize for later convenience the phases arising in this case, in the context of the Large Volume Scenario or any other decaying exponential potential for the volume modulus, sketched in Fig. 1 (see also Table 1 of [34]).  [^32] In this article we are not concerned about the mechanism driving inflation. If the volume modulus was not the inflaton, we neglect the physics that stabilises it during inflation and assume that the inflaton is stabilised after inflation. If it was the inflaton, we neglect the physics that provides a slow-roll potential in the inflationary regime. These approximations are well justified sufficiently far (in field space) from the inflationary expectation value of the modulus.
Figure 1: Schematic timeline of the scenario. Adapted from [37].
After inflation, there will be a kination phase, followed by a brief epoch of radiation/matter domination and then a radiation/matter tracker. Once the volume reaches the vacuum, it will give rise to early matter domination by oscillating around the minimum of the potential, and finally reheat the Universe by decaying to radiation. While the different transients between these epochs can partly be described analytically, we will not give all of the details here, and the interested reader is referred to Section 2 of [34]. In any case, we will also make us of numerical solutions to (3.12) to calculate GW spectra. With these assumptions in mind, a schematic timeline proceeds as follows. Kination begins at a time
$t_{0}=\frac{M_{P}}{\sqrt{3}\Lambda^{2}_{{\rm inf}}},$ (3.32)
where $\Lambda_{{\rm inf}}$ denotes the scale of inflation, i.e. $\Lambda_{{\rm inf}}=V_{0}^{1/4}e^{-\lambda\frac{\Phi_{0}}{4M_{P}}}$. During this phase, as discussed around Eq. (3.29), the string tension will evolve as $G\mu(t)\sim 1/t$. Let us now assume the existence of another fluid, with equation of state $P=(\gamma-1)\rho$ and abundance $\Omega_{\gamma}\equiv\varepsilon_{0}\ll 1$. Neglecting the potential, one finds [34]
$x(a)=\frac{1}{\sqrt{1+\varepsilon_{0}\left(\frac{a}{a}\right)^{3(2-\gamma)}}},$ (3.33)
@@
$t_{\gamma}\simeq t_{0}\varepsilon_{0}^{\frac{1}{\gamma-2}}.$ (3.34)
@@
$t\simeq t_{0}\varepsilon_{0}^{\frac{4+\gamma\left(\lambda\sqrt{\frac{2}{3}}-2\right)}{4(\gamma-2)}},$ (3.35)
@@
$t_{t}\simeq t_{0}\varepsilon_{0}^{\frac{1+\frac{\lambda}{2}+\sqrt{\frac{2}{3}}-\frac{2}{3}}{\gamma-2}}.$ (3.36)
@@
$t_{{\rm reh}}\simeq\frac{1}{\Gamma_{\Phi}},\quad\quad{\rm{with}}\quad\quad\frac{1}{48\pi}\frac{m_{\Phi}^{3}}{M_{P}^{2}}\lesssim\Gamma_{\Phi}\lesssim\frac{(\alpha_{{\rm{loop}}}\mathcal{V})^{2}}{16\pi}\frac{m_{\Phi}^{3}}{M_{P}^{2}}.$ (3.37)
@@
To close off, let us remark that the above picture is indeed schematic, with the equations being correct up to $\mathcal{O}(1)$ factor and intended to provide simple, analytical approximations to the evolution of the background useful for an intuitive understanding of the physics at play. For the actual computations of GW spectra presented in the next sections, the values of the various background cosmological parameters (scale factor, Hubble rate, string tension…) are computed numerically. In particular, we numerically solve the system (3.12) from the end of inflation to the beginning of the matter-dominated epoch, when the field starts oscillating around the minimum. From there onward, it is simpler to assume a standard radiation-matter universe, with a matter component that decays to radiation as described in 3.1.2 and with initial conditions determined by the previous evolution. Finally, this can be patched to the standard cosmological history before the onset of BBN.
@@ 4 The GW background
In this Section we follow [97] and provide analytic estimates for the amplitude and spectral index of the GW spectrum sourced by a network of cosmic superstrings, as a function of the equation of state of the background when the loops are created and when they source the dominant contribution to the spectrum. In Section 4.1 we reproduce the results of [97], generalizing them to include expressions for the amplitude as a function of $G\mu$, $\alpha$, $P$ and $\chi$ (the latter is defined in Eq. (2.14)). In Section 4.2, we generalize the computation to the case where the tension varies as a power law in time, as motivated by moduli dynamics in string theory constructions. We observe that this variation can render qualitatively different changes in the GW spectrum. In particular, it induces spectral indices that cannot be reproduced in scenarios with constant tension, and the maximum contribution to the amplitude at a given frequency is sourced at the time of formation of the loop (while with constant tension this occurs at half of its lifetime).
Let us thus compute the spectral shape of the GW background released by the loops (see [155] for a recent review). First, we assume that the loops with length $\ell$ radiate in discrete harmonics given by
$\tilde{f}_{k}=\frac{2k}{\ell}\,,\qquad k\in\mathbb{N}^{+}$ (4.1)
@@
$\Omega_{\rm{GW}}^{(1)}(f)\equiv\frac{1}{\rho_{c}}\frac{d\rho_{\rm{GW,0}}}{d\log f}=\frac{f}{\rho_{c}}\int_{t_{0}}^{t_{i}}{d\tilde{t}\,\left(\frac{a(\tilde{t})}{a_{0}}\right)^{4}\frac{dE^{(1)}_{\rm{GW}}}{d\tilde{t}}\left(\frac{a(t_{i})}{a(\tilde{t})}\right)^{3}\frac{d\ell(\tilde{t})}{df}\frac{dt_{i}}{d\ell(\tilde{t})}\frac{dn}{dt_{i}}}\,.$ (4.2)
This equation sums over all times $\tilde{t}$ the (redshifted) energy in GWs sourced from loops with length $\ell(\tilde{t})=\frac{2a(\tilde{t})}{fa_{0}}$. The density of the number of loops per unit length is written in terms of the (redshifted) initial loop distribution, taking special care of the variation of the size of the interval from $(t_{i},t_{i}+dt_{i})$ to $)(\ell(\tilde{t}),\ell(\tilde{t})+d\ell(\tilde{t}$)). In addition, $E^{(1)}_{\rm{GW}}$ denotes the fraction of energy that is released into the first mode. In general, one writes
$\frac{dE_{\rm{GW}}}{d\tilde{t}}=\sum_{k=1}^{\infty}\frac{dE^{(k)}_{\rm{GW}}}{d\tilde{t}}\,$, (4.3)
which in the language of Eq. (2.18) corresponds to
$\Gamma=\sum_{k=1}^{\infty}\Gamma^{(k)}\equiv\sum_{k=1}^{\infty}\frac{\Gamma}{k^{4/3}\zeta(4/3)}\,$, (4.4)
@@
Throughout the rest of the paper, we use $\rho_{c}^{-1}=\frac{8\pi}{3}\frac{G}{H_{0}^{2}}$ and Eqs. (2.16) and (4.4) to write the contribution from the mode $k$ to the gravitational wave spectrum as:
$\begin{split}&\Omega_{\rm{GW}}^{(k)}(f)=\frac{16\pi}{3}\frac{\Gamma^{(k)}k}{H_{0}^{2}f}\frac{\mathcal{F}_{\alpha}}{\alpha P^{2(1-\beta)+\chi}}\times\\ \int_{t_{i}}^{t_{0}}d\tilde{t}\,(G\mu(\tilde{t}))^{2}\,&\frac{C_{\rm{eff}}(t_{i})}{t_{i}^{4}}\frac{dt_{i}}{dl(\tilde{t})}\left(\frac{a(t_{i})}{a(\tilde{t})}\right)^{3}\left(\frac{a(\tilde{t})}{a_{0}}\right)^{5}\Theta\left[t_{i}-t_{\rm{osc}}\right]\Theta\left[\alpha t_{i}-\ell_{*}\right]\,.\end{split}$ (4.5)
The Heaviside theta functions ensure that the loop formation time cannot be extrapolated before the network is formed, and impose a high frequency cutoff in terms of a minimum length $\ell_{*}$ that loops can reach before other channels become more efficient (see [156] and references therein). Of course, the total spectral shape is given by the sum of these contributions, namely
$\Omega_{\rm{GW}}(f)=\sum_{k=1}^{\infty}\Omega_{\rm{GW}}^{(k)}\,$, (4.6)
@@ 4.1 Frequency dependence: constant tension
As such, one of the most interesting features of cosmic strings is that, if they are present, because of the properties of the scaling regime, their associated GW background carries information about the cosmology following their formation. In particular, the frequency dependence of the spectrum will vary depending on the equation of state of the background. In addition, a time-dependent string tension - as would be expected if the volume modulus was rolling down its potential - leads to different spectral dependence and thus provides a smoking-gun signature for scenarios as in [32, 31, 34].
The background behavior of the spectral index is well understood [97, 30], but a similar analysis with a time-dependent tension has only been carried out in a special case [87]. Let us first review the case where the string tension is constant and solely focus on the behavior of the first mode; higher modes yield a qualitatively different behavior only in the case of matter domination (the effect of higher modes has been studied analytically in [30]). To do so, we follow the logic in [97] and generalize their approach to include the dependence on the intercommutation probability and the parameter $\chi$. Let us consider two cosmological phases with
$\rho(t)=\begin{cases}\rho_{\Delta}\left(\frac{a_{\Delta}}{a(t)}\right)^{m}\,,\qquad&t<t_{\Delta}\\ \rho_{\Delta}\left(\frac{a_{\Delta}}{a(t)}\right)^{n}\,,\qquad&t\geq t_{\Delta}\end{cases}\,.$ (4.7)
In what follows we assume that the scaling regime is maintained throughout an immediate transition between the two phases in favor of an analytic understanding of the situation, but note that these issues have been studied in the past [157, 91, 19].
@@
$\left(\frac{a(t_{i})}{a(\tilde{t})}\right)^{3}=\left(\frac{a(t_{i})}{a_{\Delta}}\right)^{3}\left(\frac{a_{\Delta}}{a(\tilde{t})}\right)^{3}=\left(\frac{t_{i}}{t_{\Delta}}\right)^{6/m}\left(\frac{t_{\Delta}}{\tilde{t}}\right)^{6/n}\,.$ (4.8)
The behaviour of the spectrum with respect to $f$ is understood by studying, at a fixed $f$, which is the time $\tilde{t}$ where the loop population sources the largest contribution to the GW amplitude. Notice that the formation time carries information about the length of the loop at time $\tilde{t}$, from (2.19) we have:
$t_{i}=\frac{1}{\alpha P^{\chi}+\Gamma G\mu}\left(\frac{2k}{f}\frac{a(\tilde{t})}{a_{0}}+\Gamma G\mu\tilde{t}\right)\,,\qquad\frac{dt_{i}}{d\ell(\tilde{t})}=\frac{1}{\alpha P^{\chi}+\Gamma G\mu}.$ (4.9)
@@
$f_{\Delta}\equiv\frac{2}{\alpha P^{\chi}}\frac{1}{t_{\Delta}z_{\Delta}}\,,\qquad\frac{t_{i}}{t_{\Delta}}\simeq\frac{f_{\Delta}}{f}\left(\frac{\tilde{t}}{t_{\Delta}}\right)^{2/n}+\frac{\Gamma G\mu}{\alpha P^{\chi}}\frac{\tilde{t}}{t_{\Delta}}\,$, (4.10)
where here and in what follows we neglect $\Gamma G\mu$ over $\alpha P^{\chi}$, and $z_{\Delta}$ is the redshift at the time $t_{\Delta}$. Notice that keeping the second term in Eq. (4.10) does not contradict this statement since the relative terms could be of the same order for a choice of times and frequencies. Indeed, it is precisely this regime that turns out to be interesting for loops in cosmologies with constant string tension (and we will see how this need not be the case otherwise).
@@
$h^{2}\Omega_{\rm{GW},1}^{(1)}=A_{\Delta}\left(\frac{f}{f_{\Delta}}\right)^{3-6/m}\int_{t_{F}}^{t_{M}(f)}\frac{d\tilde{t}}{t_{\Delta}}\,\left(\frac{\tilde{t}}{t_{\Delta}}\right)^{2(6/m-4)/n+4/n}\,.$ (4.11)
$h^{2}\Omega_{\rm{GW},2}^{(1)}=A_{\Delta}\left(\frac{f}{f_{\Delta}}\right)^{-1}\left(\frac{\Gamma G\mu}{\alpha P^{\chi}}\right)^{6/m-4}\int_{t_{M}(f)}^{t_{0}}\frac{d\tilde{t}}{t_{\Delta}}\,\left(\frac{\tilde{t}}{t_{\Delta}}\right)^{6/m-4+4/n}\,.$ (4.12)
@@
$A_{\Delta}=\frac{8\pi}{3}\frac{\Gamma^{(1)}}{(H_{0}t_{0})^{2}}\frac{\mathcal{F}_{\alpha}}{\alpha P^{2(1-\beta)+\chi}}\left(G\mu_{\Delta}\right)^{2}\,C_{\rm{eff}}\left(\frac{t_{0}}{t_{\Delta}z_{\Delta}^{2}}\right)^{2}\equiv A\left(G\mu_{\Delta}\right)^{2}\,C_{\rm{eff}}\left(\frac{t_{0}}{t_{\Delta}z_{\Delta}^{2}}\right)^{2}\,$, (4.13)
where $A$ is universal and we define $G\mu_{\Delta}\equiv G\mu(t_{\Delta})$ anticipating the case where the tension varies. We note here that the effect of varying the intercommutation probability is a uniform boost in the amplitude for $\chi=0$ (as in [7]), and an additional contribution to the amplitude and a shift in the characteristic frequency $f_{\Delta}$ for $\chi\neq 0$.
In addition, we observe that the frequency dependence of the spectrum depends on the exponents in Eqs. (4.11) and (4.12), and therefore in the equation of state of the fluid governing the expansion of the Universe. For values of interest $(m,n)\in\{3,4,6\}$ the first exponent of $\tilde{t}/t_{\Delta}$ ($p_{1}$) is always larger than 1-1 while the second ($p_{2}$) is larger than 1-1 except in a matter-radiation transition (i.e. at the end of a period of early matter domination) and during matter domination. We will have more to say about such periods in Sec. 5, since the spectrum is sensitive to higher modes, but let us for the moment assume that $p_{1}>-1$ and $p_{2}<-1$. Then, the integral is dominated at a time $t_{M}$ defined as the time at which the two contributions to $t_{i}$ are equal, namely
$\frac{t_{M}}{t_{\Delta}}\equiv\left(\frac{f_{\Delta}}{f}\frac{\alpha P^{\chi}}{\Gamma G\mu}\right)^{n/(n-2)}\,.$ (4.14)
@@
$h^{2}\Omega_{\rm{GW}}^{(1)}\simeq A_{\Delta}\left(\frac{f}{f_{\Delta}}\right)^{B}\,\left(\frac{\alpha P^{\chi}}{\Gamma G\mu}\right)^{C}\,.$ (4.15)
Where B=2(mnm3n)/(m(n2))B=2(mn-m-3n)/(m(n-2)) and C=(12+mn4m)/(m(n2))C=(12+mn-4m)/(m(n-2)). There are subtleties to the analysis above that will be important later. Firstly, we must note that we have assumed that $t_{M}(f)<t_{0}$, where $t_{0}$ indicates the onset of a different phase of the Universe (eg: radiation-matter transition), or the age of the Universe today. Otherwise the dominant contribution to the spectrum at this frequency must be computed in the later phase with times $t>t_{0}$ or, in the case of the latest epoch of matter domination, it is just evaluated today. In the latter case:
$h^{2}\Omega_{\rm{GW}}^{(1)}\simeq A_{0}\left(\frac{f}{f_{0}}\right)^{3-6/m}\,.$ (4.16)
@@
$f<f_{F}\left(\frac{\alpha P^{\chi}}{\Gamma G\mu}\right)^{2/n}2^{1-2/n}\,$, (4.17)
@@
At larger frequencies one simply evaluates (4.12) at $t_{F}$, finding
$h^{2}\Omega_{\rm{GW}}^{(1)}\simeq A_{F}\left(\frac{f}{f_{K}}\right)^{-1}\left(\frac{\alpha P^{\chi}}{\Gamma G\mu}\right)^{6/m-4}\,$, (4.18)
@@
$n=3$ 1-1 1-1 11
$n=4$ 1-1 0 11
$n=6$ 1/2-1/2 $1/4$ 11
Table 1: Expected spectral index of the GW spectrum sourced by the first mode of loops that are created when the background energy density redshifts like $\rho\sim a^{-m}$ and radiate GWs with $\rho\sim a^{-n}$, according to Eq. (4.15). Negative spectral indices are altered by the effect of higher modes to 1/3-1/3.